# Transformer Layers as Painters

### v1

@@ 1 Introduction
The scale of transformer-based Large Language Models (LLMs), in the billions of parameters, makes it difficult to directly understand the models’ behaviour after training. At the same time, each layer of a pretrained transformer has an identical architecture as the other layers, with the only difference being a layer’s position in the hierarchy, and the values of the layer’s parameters Vaswani et al. (2017).
@@
    Do layers use the same representation space? (Section 3.1)
@@
    Are middle layers all performing the same function? (Section 3.3)
@@
    Does the layer order matter? (Section 3.4)
@@
    Does the order matter for some tasks more than others? (Section 3.6)
@@
    Does looping help parallelized layers? (Section 3.7)
@@
    Which variants harm performance the least? (Section 3.8)
To answer these questions we perform a series of experiments on pretrained LLMs. These include experimenting with variations on the standard transformer execution strategy, and measuring the impact of these variations on the models’ performance across a variety of benchmarks for both decoder-only (Llama) and encoder-only (BERT) models. Note that our experiments never involve finetuning or otherwise adjusting the models’ parameters (with the caveat that the GLUE evaluation standard procedure includes a finetuning step for our BERT-Large model) [^1] Code will be available at https://github.com/floatingbigcat/transformer-as-painter/..
@@ 2 Models and Benchmarks
Our experiments are primarily on two transformer models: Llama2 Touvron et al. (2023), and on BERT-Large Devlin et al. (2019). Llama2 is decoder-only. We focus on Llama2-7B, which has 7 billion parameters and 32 layers (each layer having 202 million parameters), but also include some scaling experiments with the 13B (40 layers) and 70B (80 layers) models. BERT is encoder-only with 24 layers and 340 million parameters. We used the standard pretrained checkpoints for these models. In all our experiments the models are frozen: we never modified the parameters of these models through fine-tuning or other methods, with the exception of the BERT evaluation, which includes a standard fine-tuning step.
We used standard benchmarks for both decoder-only LLMs (for Llama2) and for encoder-only LLMs (for BERT). For Llama2, we use ARC (science exam questions) Clark et al. (2018), HellaSwag (commonsense) Zellers et al. (2019), GSM8K (Math Word Problems) Cobbe et al. (2021), WinoGrande (Winograd Schema Challenge) Sakaguchi et al. (2019), and LAMBADA (word prediction) Paperno et al. (2016). This last, LAMBADA, measures perplexity and is closest to the raw token-prediction used during training. For Llama2, we include the normalized median of the benchmarks, where we scale each benchmark with 0 being the performance of random (or max-class) guessing and 1 being the performance of the full Llama2 model. For BERT, we used tasks from the GLUE benchmark Wang et al. (2018) and followed their evaluation protocol, including reporting the unnormalized average of the benchmarks. Note that standard BERT evaluation includes a fine-tuning step Devlin et al. (2019), so our BERT model has a chance to adapt to the new configuration. Therefore, we also include results from an evaluation where a head can adapt, but the model itself is frozen. These results are in Appendix A.6, and more details of the GLUE benchmark are given in Appendix A.5.
@@ 3 Experiments
The original motivation behind our experiments came from the question of whether multiple layers could be somehow be merged into a single (possibly larger) layer. (Such merging could potentially be automated Akiba et al. (2024).) We hypothesized, perhaps because of the use of residual connections during training, that the middle layers of a neural network may use a common representation space. (This is not the case for standard multi-layer perceptrons, where there is nothing to encourage a common representation or permutational consistency across layers.) The possibility of layers sharing a common representation has downstream implications for conditional computation (e.g. Pagliardini et al. (2024)) or for dynamically inserting new knowledge into pretrained transformer models.
@@ 3.1 Do layers “speak the same language”?
This experiment would suggest that the middle layers 1. share a representation space and 2. have a separate representation space from the “outer” (first and last few) layers. To further test this hypothesis, following previous work Friedman et al. (2023); Kornblith et al. (2019); Simoulin & Crabbé (2021); Godey et al. (2024); Xue et al. (2023), we measured the average cosine similarity between the activations of hidden states of different layers of our models (Llama2-7B, Llama2-13B, and BERT-Large) across our benchmarks. In Figure 3, we show that this consistency holds among all the middle layers. For example, the activation in the fourth layer from the bottom has a high similarity to the fourth layer from the top. For the 40 layers of Llama2-13B, we see that the layers form four or five distinct similarity groups: Layer 0, layers 1-3, the middle layers, then the final layer or two.
@@
Figure 3: Average cosine similarity between the hidden states of all 32 layers of Llama2-7B (Left) and all 40 layers of Llama2-13B.
@@ 4 Related Work
Kim et al. (2024) showcased that pruning entire transformers layers can reduce latency without a considerable drop in performance. This is in line with the findings in Bhojanapalli et al. (2021). Also, both the works noted that the performance drop is substantial if we drop the first few entire transformer layers. Hence there is an agreement that the first few transformers layers are crucial for performance. One implication of this observation is that many of these layers would be carrying redundant information, and this was shown by Kim et al. (2024) who removed these layers, and noticed the change in the PPL score. The authors then removed these layers in one-shot, and retrained the model with LoRA to make up for the lost performance,
One aspect where Bhojanapalli et al. (2021) and Kim et al. (2024) observations differ though is the fine-grained units. Bhojanapalli et al. (2021) observed that removing MLP layers have lesser impact on performance compared to removing an entire transformer layer, whereas Kim et al. (2024) observed that this behavior is very much dependent on the size of the models. They noted that removing individual MHA and FFN modules results in better downstream task accuracy but worse PPL compared to removing entire transformer layers when the model has more than 5B parameters. For smaller models than 5B, layer-level pruning achieves superior results. While Kim et al. (2024) did a successful job on pruning the models, the authors observed an (un)interesting side effect of the same. The pruned models perform worse when responding to factual questions or generating long responses. The authors couldn’t make up for the lost performance on these tasks even after retraining the models, suggesting that while much of the information stored in these layers was redundant, some parts of it were required for critical tasks e.g. factual Q&A.
The experiments of ShortGPT Men et al. (2024) corroborate the findings of ShortenedLlama, exploiting the redundancy in LLMs to derive a pruning technique. Denseformer Pagliardini et al. (2024) had similar findings where they found that modules even after applying DWA had cosine similarity with original transformer modules, suggesting both that there is some redundant information flow, and can be leveraged for sparsity.
@@ 5 Discussion
We leave a full explanation for why transformers are robust to our variations for future work. One possible hypothesis is that the residual connections during training are necessary for the layers to share a common representation. It’s already known that residual connections are useful to help address the vanishing gradient problem He et al. (2015), and that transformers trained without these connections perform worse than without. However, it would be interesting to rerun our variations on models without residuals, and see if our variations destroyed whatever meager gains full non-residual models achieved.
For other future work, we plan to “thaw” our models and investigate if (and how long) transformers take to adjust to the variations in the paper via fine-tuning. If these models were actually fine-tuned with their new architecture, performance would probably be even better. It’s worth noting that Parallel and Skip both have potentially lower latencies than the full model (assuming enough memory to execute the layers simultaneously). For example, the latency for the Parallel Layer for Llama2-7B for N=8 should be about half that of normal Llama2-7B. Though the aim of this paper is to better understand layers in transformer-based LLMs as opposed to introducing new models, our results suggest simple methods for easily trading accuracy for latency gains. Our results also suggest that a routing mechanism for executing frozen layers may be used for this, somewhat analogous to Switch Transformers Fedus et al. (2022).
@@ A.1 How does scaling affect hidden state similarities?
To reveal how the property of “sharing similar embeddings” evolves while increasing model size, we computed the average similarity matrices for Pythia models (Biderman et al. (2023)) ranging in size from 14M to 12B parameters, utilizing 100 data samples from the LAMBADA token prediction benchmark Paperno et al. (2016). Figure 12 shows even the tiny 70M model has a clear boundary for its beginning, middle, ending layers. Interestingly, the proportion of beginning layers decreases with an increase in model size but the number of ending layer remains consistently at a single layer across all Pythia models, regardless of size. We therefore answer our question with:
@@ A.4 Example of Error for Parallelized Layers
{boxedverbatim} Mishka spent $16.50*3=<<16.50*3=49.50>>49.50onshortsMishkaspent$22.50*3 = ¡¡22.50*3=67.50¿¿ 67.50 on pants Mishka spent
@@
{boxedverbatim} Mishka spent $16.50*3=<<16.50*3=50.50>>50.50ontheshorts.$22.50 * 3 = ¡¡22.50*3=70.50¿¿ 70.50 on the pants. $42*3=<<42*3=142>>142ontheshoes.Mishkaspent$50.50 + $70.50+$142 = ¡¡50.50+70.50+142¿¿192.50 on the clothing items. #### 192.50
@@ A.5 BERT benchmarks
Below are the benchmarks we used from GLUE Wang et al. (2018):

### v2

@@ 1 Introduction
The scale of transformer-based Large Language Models (LLMs), in the billions of parameters, makes it difficult to directly understand the models’ behaviour after training. At the same time, each layer of a pretrained transformer has an identical architecture as the other layers, with the only difference being a layer’s position in the hierarchy, and the values of the layer’s parameters (Vaswani et al. 2017).
@@
    Do layers use the same representation space? (§3.1)
@@
    Are middle layers all doing the same function? (§3.3)
@@
    Does the layer order matter? (§3.4)
@@
    Does order matter for some tasks more than others? (§3.6)
@@
    Does looping help parallelized layers? (§3.7)
@@
    Which variants harm performance the least? (§3.8)
To answer these questions we perform a series of experiments on pretrained LLMs. These include experimenting with variations on the standard transformer execution strategy, and measuring the impact of these variations on the models’ performance across a variety of benchmarks for both decoder-only (Llama) and encoder-only (BERT) models. Note that our experiments never involve finetuning or otherwise adjusting the models’ parameters (with the caveat that the GLUE evaluation standard procedure includes a finetuning step for our BERT-Large model) [^1] Our code is available at https://anonymous.4open.science/r/transformer-as-painter-CF60.
@@ 2 Models and Benchmarks
Our experiments are primarily on two transformer models: Llama2 (Touvron et al. 2023), and on BERT-Large (Devlin et al. 2019). (However, we also include results for Mistral-7B (Jiang et al. 2023) and Pythia-6.9B (Biderman et al. 2023a) in Appendix A.7 that support the generalization of our results.) Llama2 is decoder-only. We focus on Llama2-7B, which has 7 billion parameters and 32 layers (each layer having 202 million parameters), but also include some scaling experiments with the 13B (40 layers) and 70B (80 layers) models. BERT is encoder-only with 24 layers and 340 million parameters. We used the standard pretrained checkpoints for these models. In all our experiments the models are frozen: we never modified the parameters of these models through fine-tuning or other methods, with the exception of the BERT evaluation, which includes a standard fine-tuning step.
We used standard benchmarks for both decoder-only LLMs (for Llama2) and for encoder-only LLMs (for BERT). For Llama2, we use ARC (science exam questions) (Clark et al. 2018), HellaSwag (commonsense) (Zellers et al. 2019), GSM8K (Math Word Problems) (Cobbe et al. 2021), WinoGrande (Winograd Schema Challenge) (Sakaguchi et al. 2019), and LAMBADA (word prediction) (Paperno et al. 2016). This last, LAMBADA, measures perplexity and is closest to the raw token-prediction used during training. For Llama2, we include the normalized median of the benchmarks, where we scale each benchmark with 0 being the performance of random (or max-class) guessing and 1 being the performance of the full Llama2 model. For BERT, we used tasks from the GLUE benchmark (Wang et al. 2018) and followed their evaluation protocol, including reporting the unnormalized average of the benchmarks. Note that standard BERT evaluation includes a fine-tuning step (Devlin et al. 2019), so our BERT model has a chance to adapt to the new configuration. Therefore, we also include results from an evaluation where an additional output layer can adapt, but the model itself is frozen. These results are in Appendix A.6, and more details of the GLUE benchmark are given in Appendix A.5.
@@ 3 Experiments
The original motivation behind our experiments came from the question of whether multiple layers could be somehow be merged into a single (possibly larger) layer. (Such merging could potentially be automated (Akiba et al. 2024).) We hypothesized, perhaps because of the use of residual connections during training, that the middle layers of a neural network may use a common representation space. (This is not the case for standard multi-layer perceptrons, where there is nothing to encourage a common representation or permutational consistency across layers.) The possibility of layers sharing a common representation has downstream implications for conditional computation (e.g. (Pagliardini et al. 2024)) or for dynamically inserting new knowledge into pretrained transformer models.
@@ 3.1 Do layers “speak the same language”?
This experiment would suggest that the middle layers 1. share a representation space and 2. have a separate representation space from the “outer” (first and last few) layers. To further test this hypothesis, following previous work (Friedman et al. 2023; Kornblith et al. 2019; Simoulin and Crabbé 2021; Godey, Éric de la Clergerie, and Sagot 2024; Xue et al. 2023), we measured the average cosine similarity between the activations of hidden states of different layers of our models (Llama2-7B, Llama2-13B, and BERT-Large) across our benchmarks. In Figure 3, we show that this consistency holds among all the middle layers. For example, the activation in the fourth layer from the bottom has a high similarity to the fourth layer from the top. For the 40 layers of Llama2-13B, we see that the layers form four or five distinct similarity groups: Layer 0, layers 1-3, the middle layers, then the final layer or two.
@@
Figure 3: Avg. cosine similarity between the hidden states of all 32 layers of Llama2-7B (Left) and all 40 layers of Llama2-13B.
@@ 4 Related Work
(Kim et al. 2024) showcased that pruning entire transformers layers can reduce latency without a considerable drop in performance. This is in line with the findings in (Bhojanapalli et al. 2021). Also, both the works noted that the performance drop is substantial if we drop the first few entire transformer layers. Hence there is an agreement that the first few transformers layers are crucial for performance. One implication of this observation is that many of these layers would be carrying redundant information, and this was shown by (Kim et al. 2024) who removed these layers, and noticed the change in the PPL score. The authors then removed these layers in one-shot, and retrained the model with LoRA to make up for the lost performance,
One aspect where (Bhojanapalli et al. 2021) and (Kim et al. 2024) observations differ though is the fine-grained units. (Bhojanapalli et al. 2021) observed that removing MLP layers have lesser impact on performance compared to removing an entire transformer layer, whereas (Kim et al. 2024) observed that this behavior is very much dependent on the size of the models. They noted that removing individual MHA and FFN modules results in better downstream task accuracy but worse PPL compared to removing entire transformer layers when the model has more than 5B parameters. For smaller models than 5B, layer-level pruning achieves superior results. While (Kim et al. 2024) did a successful job on pruning the models, the authors observed an (un)interesting side effect of the same. The pruned models perform worse when responding to factual questions or generating long responses. The authors couldn’t make up for the lost performance on these tasks even after retraining the models, suggesting that while much of the information stored in these layers was redundant, some parts of it were required for critical tasks e.g. factual Q&A.
The experiments of ShortGPT (Men et al. 2024) corroborate the findings of ShortenedLlama, exploiting the redundancy in LLMs to derive a pruning technique. Denseformer (Pagliardini et al. 2024) had similar findings where they found that modules even after applying DWA had cosine similarity with original transformer modules, suggesting both that there is some redundant information flow, and that this can be leveraged for sparsity.
More recently, (Freiberger et al. 2024) explores layer shuffling during training to enhance robustness of the models, while (Dutta, Gupta, and Agarwal 2024) proposes an algorithm that can be used for efficient pruning of transformers. (Lad, Gurnee, and Tegmark 2024) explores the robustness of transformer-based LLMs by deleting or swapping layers. (Zou et al. 2024) focuses on efficient inference by splitting layers in groups, running them in parallel or bypassing them. On a similar note, (Flynn et al. 2024) focuses on pruning transformers in different ways (entire attention blocks, ffn, etc.). Our work is more closely related to (Lad, Gurnee, and Tegmark 2024) and (Flynn et al. 2024) where the ablations involve frozen models. We present a super set of such ablations for the frozen transformer models.
@@ 5 Discussion
We leave a full explanation for why transformers are robust to our variations for future work. One possible hypothesis is that the residual connections during training are necessary for the layers to share a common representation. It’s already known that residual connections are useful to help address the vanishing gradient problem (He et al. 2015), and that transformers trained without these connections perform worse than without. However, it would be interesting to rerun our variations on models without residuals, and see if our variations destroyed whatever meager gains full non-residual models achieved.
For other future work, we plan to “thaw” our models and investigate if (and how long) transformers take to adjust to the variations in the paper via fine-tuning. If these models were actually fine-tuned with their new architecture, performance would probably be even better. It’s worth noting that Parallel and Skip both have potentially lower latencies than the full model (assuming enough memory to execute the layers simultaneously). For example, the latency for the Parallel Layer for Llama2-7B for N=8 should be about half that of normal Llama2-7B. Though the aim of this paper is to better understand layers in transformer-based LLMs as opposed to introducing new models, our results suggest simple methods for easily trading accuracy for latency gains. Our results also suggest that a routing mechanism for executing frozen layers may be used for this, somewhat analogous to Switch Transformers (Fedus, Zoph, and Shazeer 2022).
@@ A.1 How does scaling affect hidden state similarities?
To reveal how the property of “sharing similar embeddings” evolves while increasing model size, we computed the average similarity matrices for Pythia models ((Biderman et al. 2023b)) ranging in size from 14M to 12B parameters, utilizing 100 data samples from the LAMBADA token prediction benchmark (Paperno et al. 2016). Figure 12 shows even the tiny 70M model has a clear boundary for its beginning, middle, ending layers. Interestingly, the proportion of beginning layers decreases with an increase in model size but the number of ending layer remains consistently at a single layer across all Pythia models, regardless of size. We therefore answer our question with:
@@ A.4 Example of Error for Parallelized Layers
{boxedverbatim} Mishka spent $16.50*3=<<16.50*3=49.50>>49.50onshortsMishkaspent$22.50*3 = <<22.50*3=67.50>> 67.50 on pants Mishka spent
@@
{boxedverbatim} Mishka spent $16.50*3=<<16.50*3=50.50>>50.50ontheshorts.$22.50 * 3 = <<22.50*3=70.50>> 70.50 on the pants. $42*3=<<42*3=142>>142ontheshoes.Mishkaspent$50.50 + $70.50+$142 = <<50.50+70.50+142>>192.50 on the clothing items. #### 192.50
@@ A.5 BERT benchmarks
Below are the benchmarks we used from GLUE (Wang et al. 2018):
@@ A.6 Results for Frozen BERT
A.7 Do the results hold for Mistral and Pythia?
To further investigate the generalization of our findings, we include summaries our results for Mistral-7B (Jiang et al. 2023) and Pythia-6.9B. Mistral-7B is a decoder-only model whose architecture is largely similar to Llama2-7B, but with some important modifications such as a sliding window attention and grouped-query attention, and, of course, different weight values. As the name suggests, Pythia-6.9B (Biderman et al. 2023a) is a 6.9 billion parameter decoder model, trained on public data, with the same architecture as the Open Pre-trained Transformer OPT-6.7B (Zhang et al. 2022).
As can be seen in the figure below Mistral follows a remarkably similar trend as Llama2-7B, which isn’t surprising given the similarity of the architectures.
However, in the figure below, we see that Pythia is less robust to modifications, especially modification of the Layer order compared to Mistral and Llama2-7B.